If Cos x/Cos y=2 and Cos(x-y)=√3/2 then tan y is  how much.this qjequest is from trigonometry

Grade:12th pass

1 Answers

Arun
5 years ago

cos(x-y)=sqrt3/2

cos(pi/6)=sqrt3/2

x-y = pi/6

x=y+pi/6

replace x by y+pi/6 in the other equation given.

cos x/cos y =2

cos (y+pi/6)=2*cos y

(cos y)(cos pi/6)-(sin y)(sin pi/6) = 2*cos y

dividing all by cos y get cos pi/6 - (tan y)*1/2 = 2

sqrt 3/2 - (tan y) *1/2 = 2

tan y = sqrt 3 -4